How much heat is required to totally melt 6 kg of aluminum which is at 30 °C.


1
Expert's answer
2021-05-10T02:42:37-0400


aluminium has a specific heat of 0.9KJ/KgK, a melting point of 660 degrees Celsius

Formula adopted

dQ= m c dt

m=6kg

c=0.9KJ/KgK

dt=660-30=630 deg celcius=630+273K=903K

dQ= 6* 0.90*903=4876.2kJ